How is the rate of a chemical reaction expressed? Explain giving a suitable
example.
1
Expert's answer
2019-10-23T07:01:04-0400

Rate of chemical reaction is the measure of how fast these changes are taking place.

 It is expressed in terms of either the concentration (amount per unit volume) of a product that is formed in a unit of time or the concentration of a reactant that is consumed in a unit of time.

For example: Let us assume an elementary reaction such that:

"A+2B\to C"

"r=-\frac{dA}{dt}=-\frac{1}{2}\frac{dB}{dt}=\frac{dC}{dt}=k[A][B]^2"